# GPU-accelerated Evolutionary Multiobjective Optimization Using Tensorized RVEA

### v1

@@ Abstract.
copyright: acmlicenseddoi: 10.1145/nnnnnnn.nnnnnnnisbn: 978-x-xxxx-xxxx-x/YY/MMconference: The Genetic and Evolutionary Computation Conference 2024; July 14–18, 2024; Melbourne, Australiajournalyear: 2024ccs: Theory of computation Evolutionary algorithmsccs: Applied computing Multi-criterion optimization and decision-making
@@ 1. Introduction
GPUs exhibit advanced computational capabilities in many fields (Owens et al., 2008), primarily due to their outstanding parallel processing abilities. In particular, the JAX (Bradbury et al., 2018) framework has been instrumental in advancing GPU-accelerated libraries such as EvoJAX (Tang et al., 2022), evosax (Lange, 2023), and EvoX (Huang et al., 2023), which enhance evolutionary computation and offer performance improvements over traditional CPU-based methods. These developments are crucial for addressing computational challenges in EMO, especially in LSMOPs. However, the application of GPU acceleration in EMO algorithms has predominantly been concentrated on enhancing the NSGA-II, with other EAs receiving far less focus (Gupta and Tan, 2015; Wong, 2009).
@@ 3. Tensorization of RVEA
In the proposed TensorRVEA, a comprehensive implementation of tensorization is achieved through two principal components: the tensorization of data structures and the tensorization of operators. While retaining the overarching framework of the original RVEA, TensorRVEA introduces significant advancements in data handling and operational efficiency. Additionally, TensorRVEA is implemented on EvoX (Huang et al., 2023). This integration ensures that the enhanced capabilities of tensorization are seamlessly incorporated within the original RVEA structure, thereby optimizing the algorithm’s performance in complex optimization scenarios. The main notations used in this paper are summarized in Table 1.
@@ 4. Experiments
This study encompasses four comprehensive experiments to rigorously evaluate TensorRVEA: acceleration performance, numerical benchmarks, multiobjective neuroevolution, and experiments demonstrating its extensibility by altering several reproduction operators. For fair comparisons, all experiments were consistently conducted on the EvoX platform (Huang et al., 2023), utilizing a system equipped with an AMD EPYC 7543 8-Core Processor server and an NVIDIA® GeForce RTX 4090 GPU. Detailed experimental setups are provided in Appendix C.
@@ 4.3. Multiobjective Neuroevolution
Figure 7. Performance comparison of TensorRVEA integrated with CSO, DE, GA, PSO and Random (denoted as Rand) operators on MoHopper-m2 and MoHopper-m3 problems. Note: larger values indicate better performances.
@@
As shown in Figure 6, TensorRVEA shows the best performance on all problems. The final average HV and average EU scores of TensorRVEA in the four environments surpass those of NSGA-II and random search. However, it is important to note that the confidence intervals for TensorRVEA tend to increase during the later phases, suggesting a potential variability in its performance outcomes. In contrast, the confidence intervals for both NSGA-II and random search remain comparatively narrow. Interestingly, the HV performance of NSGA-II in the MoHalfCheetah and MoSwimmer environments does not match the efficacy of random search. This observation points to the varying effectiveness of these algorithms across different problem settings. In the MoHopper-m3 environment, NSGA-II and RS show similar performance in the HV and the EU metrics. This resemblance could be attributed to the dominance resistance phenomenon and a swift increase in nondominated solutions. TensorRVEA consistently yields high-quality and diverse policies across all tested environments, underlining its versatility. This superior performance is attributable to its selection mechanism based on reference vectors. Unlike the non-dominated sorting method, this method facilitates evolution in the direction of the reference vectors, enabling it to effectively escape local optima. Notably, diversity is a crucial factor in robot control tasks, catering to varied requirements such as high-speed operation at the expense of energy efficiency or the converse. In the MoSwimmer environment, NSGA-II produces more evenly distributed solutions, yet there remains room for improvement in solution quality. Meanwhile, the solutions generated by random search tend to lose diversity, with multiple solutions emerging at specific performance levels.
@@ 4.4. Extensibility
In this experiment, the extensibility of TensorRVEA was tested by employing a variety of reproduction operators. Within the framework of TensorRVEA, we integrated several representative reproduction operators, including Genetic Algorithm (GA) (Deb et al., 1996a), Particle Swarm Optimization (PSO) (Eberhart and Kennedy, 1995), Differential Evolution (DE) (Storn and Price, 1997), Competitive Swarm Optimizer (CSO) (Cheng and Jin, 2015), as well as the Random generation serving as the baseline for comparison. Here, CSO is picked, as it is tailored for large-scale optimization, a required feature for neuroevolution. The population size was fixed at 10,000, and each algorithm was executed over 100 generations in both the 2-objective and 3-objective MoHopper environments. The performance of these variations was evaluated in terms of HV and EU.
@@ 5. Conclusion
For advancing evolutionary multiobjective optimization (EMO) towards addressing massive-scale and many-objective problems, our work demonstrates the unique advancements of tensorization and GPU acceleration, using TensorRVEA as the instance. The robustness and efficiency of TensorRVEA is assessed on a diverse range of optimization challenges, including traditional numerical benchmark and high-dimensional neuroevolution challenges for robot control tasks. The experimental results highlight TensorRVEA’s notable computational speed and enhanced performance, which is attributed to its fully tensorized design and GPU-accelerated implementation. Furthermore, the algorithm’s adaptability in integrating various reproduction operators further underscores its broad extensibility. Looking ahead, we will further explore the tensorization of general EMO algorithms, probably towards the development of tailored algorithm frameworks and operators for GPU acceleration.
@@ C.2. Numerical benchmarks
Table 5. Experimental setup for assessing model performance on DTLZ problems
@@ C.3. Multiobjective Neuroevolution
Figure 9. Robot control tasks involved in the experiment for multiobjective neuroevolution.
@@ C.3.2. Experimental Setup
Stop Time MoHalfCheetah: 15000s, MoHopper-m2: 400s,
MoSwimmer: 800s, MoHopper-m3: 400s
@@
When computing the Hypervolume (HV), as delineated in Eq. 13, the reference point is established by taking the worst values from the Pareto solutions across all iterations of the TensorRVEA, NSGA-II, and the random search algorithms. If the worst values of the objectives, specifically forward reward and height, fall below zero, the reference points for these objectives are established at zero. Table 7 illustrates the reference points for each environment used in the computations.
@@ C.4. Extensibility
The experimental setups of extensibility experiments are shown in Table 8 and Table 9 illustrates the reference points for each environment used in the computations. The reference point is established by taking the worst values from the Pareto solutions across all iterations of the four variants of the TensorRVEA algorithm. If the worst values of the objectives, specifically forward reward and height, fall below zero, the reference points for these objectives are established at zero.
@@
Reproduction Operators GA, PSO, DE, CSO, Random Generation (baseline)

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2024; Melbourne, VIC, Australiabooktitle: Genetic and Evolutionary Computation Conference (GECCO ’24), July 14–18, 2024, Melbourne, VIC, Australiadoi: 10.1145/3638529.3654223isbn: 979-8-4007-0494-9/24/07ccs: Theory of computation Evolutionary algorithmsccs: Applied computing Multi-criterion optimization and decision-making
@@ 1. Introduction
GPUs exhibit advanced computational capabilities in many fields (Owens et al., 2008), primarily due to their outstanding parallel processing abilities. In particular, the JAX (Bradbury et al., 2018) framework has been instrumental in advancing GPU-accelerated libraries such as EvoJAX (Tang et al., 2022), evosax (Lange, 2023), and EvoX (Huang et al., 2024), which enhance evolutionary computation and offer performance improvements over traditional CPU-based methods. These developments are crucial for addressing computational challenges in EMO, especially in LSMOPs. However, the application of GPU acceleration in EMO algorithms has predominantly been concentrated on enhancing the NSGA-II, with other EAs receiving far less focus (Gupta and Tan, 2015; Wong, 2009).
@@ 3. Tensorization of RVEA
In the proposed TensorRVEA, a comprehensive implementation of tensorization is achieved through two principal components: the tensorization of data structures and the tensorization of operators. While retaining the overarching framework of the original RVEA, TensorRVEA introduces significant advancements in data handling and operational efficiency. Additionally, TensorRVEA is implemented on EvoX (Huang et al., 2024). This integration ensures that the enhanced capabilities of tensorization are seamlessly incorporated within the original RVEA structure, thereby optimizing the algorithm’s performance in complex optimization scenarios. The main notations used in this paper are summarized in Table 1.
@@ 4. Experiments
This study encompasses four comprehensive experiments to rigorously evaluate TensorRVEA: acceleration performance, numerical benchmarks, multiobjective neuroevolution, and experiments demonstrating its extensibility by altering several reproduction operators. For fair comparisons, all experiments were consistently conducted on the EvoX platform (Huang et al., 2024), utilizing a system equipped with an AMD EPYC 7543 8-Core Processor server and an NVIDIA® GeForce RTX 4090 GPU. Detailed experimental setups are provided in Appendix C.
@@ 4.3. Multiobjective Neuroevolution
Figure 7. Performance comparison of TensorRVEA integrated with CSO, DE, GA, PSO and Random reproduction (denoted as Rand) operators on MoHopper-m2 and MoHopper-m3 problems. Note: larger values indicate better performances.
@@
As shown in Figure 6, TensorRVEA shows the best performance on all problems. The final average HV and average EU scores of TensorRVEA in the four environments surpass those of NSGA-II and random search. However, it is important to note that the confidence intervals for TensorRVEA tend to increase during the later phases, suggesting a potential variability in its performance outcomes. In contrast, the confidence intervals for both NSGA-II and random search remain comparatively narrow. Interestingly, the HV performance of NSGA-II in the MoHalfCheetah and MoSwimmer environments does not match the efficacy of random search. This observation points to the varying effectiveness of these algorithms across different problem settings. In the MoHopper-m3 environment, NSGA-II and RS show similar performance in the HV and the EU metrics. This resemblance could be attributed to the dominance resistance phenomenon and a swift increase in nondominated solutions. TensorRVEA consistently yields high-quality and diverse policies across all tested environments, underlining its versatility. This superior performance is attributable to its selection mechanism based on reference vectors. Unlike the non-dominated sorting method, this method facilitates evolution in the direction of the reference vectors, enabling it to effectively escape local optima. Notably, diversity is a crucial factor in robotic control tasks, catering to varied requirements such as high-speed operation at the expense of energy efficiency or the converse. In the MoSwimmer environment, NSGA-II produces more evenly distributed solutions, yet there remains room for improvement in solution quality. Meanwhile, the solutions generated by random search tend to lose diversity, with multiple solutions emerging at specific performance levels.
@@ 4.4. Extensibility
In this experiment, the extensibility of TensorRVEA was tested by employing a variety of reproduction operators. Within the framework of TensorRVEA, we integrated several representative reproduction operators, including Genetic Algorithm (GA) (Deb et al., 1996a), Particle Swarm Optimization (PSO) (Eberhart and Kennedy, 1995), Differential Evolution (DE) (Storn and Price, 1997), Competitive Swarm Optimizer (CSO) (Cheng and Jin, 2015), as well as the Random reproduction serving as the baseline for comparison. Here, CSO is picked, as it is tailored for large-scale optimization, a required feature for neuroevolution. The population size was fixed at 10,000, and each algorithm was executed over 100 generations in both the 2-objective and 3-objective MoHopper environments. The performance of these variations was evaluated in terms of HV and EU.
@@ 5. Conclusion
For advancing evolutionary multiobjective optimization (EMO) towards addressing massive-scale and many-objective problems, our work demonstrates the unique advancements of tensorization and GPU acceleration, using TensorRVEA as the instance. The robustness and efficiency of TensorRVEA is assessed on a diverse range of optimization challenges, including traditional numerical benchmark and high-dimensional neuroevolution challenges for robotic control tasks. The experimental results highlight TensorRVEA’s notable computational speed and enhanced performance, which is attributed to its fully tensorized design and GPU-accelerated implementation. Furthermore, the algorithm’s adaptability in integrating various reproduction operators further underscores its broad extensibility. Looking ahead, we will further explore the tensorization of general EMO algorithms, probably towards the development of tailored algorithm frameworks and operators for GPU acceleration.
@@ C.2. Numerical benchmarks
Table 5. Experimental setup for conducting numerical benchmarks on DTLZ problems
@@ C.3. Multiobjective Neuroevolution
Figure 9. Robotic control tasks involved in the experiment for multiobjective neuroevolution.
@@ C.3.2. Experimental Setup
[…]
@@
In the computation of the Hypervolume (HV), as detailed in Eq. 13, the reference point is established by selecting the minimum objective values from the Pareto-optimal solutions obtained in all iterations of the TensorRVEA, NSGA-II, and random search algorithms. If the minimum values of the objectives, particularly forward reward and height, are less than zero, the reference points for these objectives are set at zero. Table 7 details the reference points for each environment used in the HV computations.
@@ C.4. Extensibility
The experimental setups of the extensibility experiments are detailed in Table 8, while Table 9 presents the reference points established for each environment in the computations. These reference points are derived by selecting the minimal objective values from the Pareto-optimal solutions, accumulated over all iterations of the five variants of the TensorRVEA algorithm. In cases where the minimal values of certain objectives, such as forward reward and height, fall below zero, the reference points for these objectives are set to zero.
@@
Reproduction Operators GA, PSO, DE, CSO, Random reproduction (baseline)